rachel works as a carpenter and earn $ 22 per hour she has a dentist appointment which will take 3 hours .she must take a texi for $15 each way to and from the appointment and the dentist charges $120 for the checkup.calculate the opportunity cost for rachel togo to the dentist.
1
Expert's answer
2018-04-23T09:20:07-0400
22*3+15*2+120=66+30+120=$216